OpenStudy (naga447):

Graphing help???? Please!!! Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). y = -one thirdx - 6 y = -one thirdx - 14 y = 3x - 6 y = 3x - 14

OpenStudy (mikurout):

1st find the slope of given line.

OpenStudy (mikurout):

Slope of given line will be equal to the slope of require line.

OpenStudy (naga447):

\[y1 - y2 \over x2 - x1\]

OpenStudy (mikurout):

Then put the value of slope in general line equation y= mx +c

OpenStudy (naga447):

I use this to find the slope right?

OpenStudy (mikurout):

this formula is for finding the slope in between two points.

OpenStudy (texaschic101):

you can just put 3x - y = -3 into y = mx + b form, and the slope is in the m position

OpenStudy (mikurout):

there is another simple technique , that is use the general line equation y = mx +c here m is the slope.

OpenStudy (naga447):

okay but what does c stand for in you equation?

OpenStudy (texaschic101):

the last letter, whether it is c or b is the y intercept

OpenStudy (mikurout):

c is the intercept of the line with the y- axis.

OpenStudy (mikurout):

are you getting this thing?

OpenStudy (naga447):

Yeah I think that I understand the equation that I need to use would look like this then right? y = -3 * 4 +-2

OpenStudy (mikurout):

no , tell me what is value of slope in the given line?

OpenStudy (texaschic101):

3x - y = -3 --- put this in y = mx + b form 3x + 3 = y -- rearrange y = 3x + 3 when you put your equation in y = mx + b form, the m is your slope. so in y = 3x + 3, the slope is 3. Do you understand so far ?

OpenStudy (mikurout):

@naga447 .follow the procedure given by @texaschic101 .

OpenStudy (naga447):

okay but did you mean to switch the -3 to a positive 3?

OpenStudy (texaschic101):

let me do this step by step.. 3x - y = -3 (add y to both sides) 3x - y + y = y - 3 (combine like terms) 3x = y - 3 (add 3 to both sides) 3x + 3 = y - 3 + 3 (combine like terms) 3x + 3 = y rearranged is : y = 3x + 3 understand ? And the number in the m position is the slope. have I lost you ?

OpenStudy (naga447):

Yeah I believe that I get it the slope is 3 correct?

OpenStudy (texaschic101):

yes..the slope is 3. And since we are looking for a parallel line, the slope will remain the same. If we were looking for a perpendicular line, it would be different. So now we have to use a slope of 3. And we have (4,-2) as our points. Now we can sub this back into y = mx + b slope(m) = 3 (4,-2)...x = 4 and y = -2 and since we know the slope is 3, we will be looking for b, the y intercept. so we sub y = mx + b -2 = 3(4) + b can you finish and solve for b ?

OpenStudy (naga447):

-2 = 3 * 4 + b -2 / 4 = 3 + b -0.5 = 3 + b 3 - -0.5 = b 3.5 = b?

OpenStudy (texaschic101):

-2 = 3(4) + b -2 = 12 + b --- subtract 12 from both sides -2 - 12 = 12 - 12 + b -- combine like terms -14 = b what you do on one side of the equal sign, has to be done on the other side. I had to isolate b, and to do this, I had to subtract 12 from both sides so that the 12 on the right side gets eliminated. Do you understand so far ?

OpenStudy (texaschic101):

so we have found our slope to be 3 and our y intercept to be -14. And in y = mx + b form, the m is the slope and the b is the y intercept. So the parallel line will be : y = 3x - 14 any questions ?

OpenStudy (texaschic101):

if you have any questions, please ask. I am here to help you :)

OpenStudy (naga447):

That makes a lot more sense then what I was trying to do! Thank you very much for helping me out and putting up with my stupidity!